OpenStudy (anonymous):

The Sumner family has 5 children. Patty is 7 years younger than her older brother, Adam. Tracy and Adam are twins. Janelle is half as old as Patty. Brian, who is 13, is 6 years older than Janelle. How old is Adam?

OpenStudy (anonymous):

P+7=A T=A 2J=P B=13=J+6

OpenStudy (anonymous):

J=7 (B=13=J+6) P=14 (2J=P) A=21 (P+7=A)

OpenStudy (anonymous):

Each letter denotes initials of names